Run Planning Meeting Notes, Mar 29-Apr 4, 2018

From GlueXWiki
Jump to: navigation, search

<- Back to Run Coordination Meetings:Fall2017 Spring2018 Run

Thursday, March 28 at 08:45

  1. Notes from the previous 24 hours:
    • RF separator tested out ok, aimed to restore beam for swing shift but delayed due to CHL issues
    • Beam arrived just before midnight, convergence not good, owl shift spent tuning
    • 0800 MCC meeting notes - Todd retuning; planning machine work 4h Mon, 8h Tues MCC whiteboard
  2. Issues:
    • Shift-takers should at least log any acknowledged alarms so that they are not forgotten
    • Is there a how-to page for the BTA sheet? (load from epics, edit etc & guidelines on ABU/BANU)
  3. Reports:
  4. Run Plan
    • Re-establish Photon beam
    • 1h production to check beam quality & detectors
    • Diamond spot adjustment (Hovanes, needs to start before noon or try again tomorrow)
    • CDC Background Studies (Lubomir et al.)
    • Converter thickness study (Alex S.)
    • SSP Study (Alex S., postpone to early next week)
    • Production running
  5. AOB

Friday, March 29 at 08:45

  1. Notes from the previous 24 hours:
      • Todd retuned Hall D beamline Thursday morning, considerable improvement
      • Hovanes set diamond positions but didn't have time for spot adjustment, #1 for today
      • CDC background studies completed ?
      • Beam became less reliable throughout swing shift, improved after midnight
    • 0800 MCC meeting notes - beam studies 0830-1230 Monday
    • MCC whiteboard
  2. Issues:
    • Some rootspy hiccups, hopefully fixed
    • Beam steering locks on 5C11B were off, there should be locks on 5C11B and the active collimator
    • Tagger quadrupole current had loaded incorrect default value during tuning! MCC will correct the default and add alarms.
    • BTA guidelines
  3. Reports:
  4. Run Plan
    • Schematic
    • Diamond spot adjustment (Hovanes, 2h) - done
    • Converter thickness study (Alex S., 2h) - done, needs to check data quality
    • CDC high thresholds (Beni, 2h) - done
    • Production running through weekend, with possible 0.5h trigger work from Alex S.
    • Over the weekend Richard will do prep work for the TAC run on Wed.
  5. Next meeting: MONDAY 0845
  6. Thanks and best wishes to Tom Carstens who retires today

Monday, April 2 at 08:45

  1. Notes from the previous 72 hours:
    • Friday morning Hovanes repositioned the diamond TagM; Sasha took converter thickness data
    • Friday pm started CDC high threshold runs but lost beam due to Hall C's boiling study
    • Friday eve Accelerator recovery from boiling study was slow. Weather event. Beam position inconsistent, MCC having trouble with the locks
    • Saturday early morning (owl) beam position became more stable
    • Saturday morning RF trip, no beam all morning. Mentioned trouble w beam position/locks not working to PD during morning check-in.
    • Saturday pm CDC threshold runs done; down for 3h w LCW fault; resumed production ~ 8pm
    • Saturday late swing and owl beam position good, steady overnight production
    • Sunday morning good production running.
    • Sunday swing first half of the shift saw RF cavity problems, then improved
    • MCC meeting notes: beam studies 4->6h, probably follow w Hall C boiling studies. Tomorrow 8h RF recovery.
    • MCC whiteboard
  2. Issues:
    • Gas shed key is absent but we know where it is
    • Ops logbook entry on locks
    • Underfocused - ask Todd for retune on Monday pm?
    • Occasional rootspy seg faults. Digihits occupancy for PSC looks strange in monitoring plots. Sasha says maybe rootspy issue?
    • Best time for harp scans is 0830+
    • Easter egg hunt (find the popup) - Hovanes will look at this after the run period ends
  3. Reports:
  4. Run Plan
    • Schematic
    • Beam studies 0830+ probably followed by Hall C's boiling studies.
    • Access requests for Monday: Lubomir. Hovanes (remove tungsten, add Pb converter to beam profiler as prep for TAC run). Elton/Sasha (for BCAL fadc). Scot (solenoid)
      • Need Radcon to survey collimator cave
    • Should have beam by 6pm. Asked PD to request a retune from Todd and/or MCC. Then do harp scan.
    • TAC run prep - add test w V wire for this evening. (Richard).
  5. AOB

Tuesday, April 3 at 08:45

  1. Notes from the previous 24 hours:
    • Access during Monday day-shift - Lubomir checked noisy channel, Hovanes (remove tungsten, add Pb converter to beam profiler), Elton (BCAL signal is fine, fadc is missing bits, this is not critical), Scot corrected solenoid probe orientation.
    • Recovered beam from Hall C studies late in swing. Retune - convergence slightly better Rates after convergence adjustment
    • Steady production overnight
    • TAC run prep - active collimator test (no V wire) (Richard) - done
    • MCC meeting notes - Todd to look at convergence tomorrow morning with aim of starting TAC run at 0930
    • MCC whiteboard
  2. Issues:
  3. Reports:
  4. Run Plan
    • Schematic
    • 8h RF recovery
    • Harp scan - done
    • TOF threshold study (Beni, evening shift, 2h) - done
    • 20 min raw mode run - done
    • Access requests for Tuesday: Colin (replace FCAL bases), Sasha (replace fadc), Scot
  5. AOB

Wednesday, April 4 at 08:45

  1. Notes from the previous 24 hours:
    • Access during Tuesday day-shift - FCAL bases & BCAL fadc replaced; solenoid work
    • Recovered beam early in swing. harp scan, TOF studies, raw mode done.
    • Late swing & owl beam off ~ 3-4h, then production
    • MCC meeting notes
    • MCC whiteboard
  2. Issues:
    • FDC timing - 2 issues, config & calibration. Shift workers need to check FDC wire timing histo. If the downloaded configuration fails readback tests, the DAQ does not start. Call the DAQ number. Sasha says if the configuration somehow downloaded incorrectly and times shifts are >= 100ns, reset run. If it is a calibration issue, notify Sean.
    • PS coincidence rates - need to monitor these & restart if frozen. Should be approx 2600Hz w diamond, 1600Hz amo PS reset instructions Need to add "hit start" to instructions.
    • epics vars swapped Q from Richard
    • Beam profiler left in yesterday
    • Rootspy restart process changed - see whiteboard
    • Phone numbers - status of on call list posted on far window vs cellphone list on whiteboard, David will clean up.
    • BEM
  3. Reports:
  4. Run Plan
    • Schematic
    • Todd to look at convergence this morning
    • TAC run 0930?? postponed.
    • Production
  5. AOB
    • Request for higher current run(s) from Sergey & Lubomir, to be discussed for later in the week.
    • Request for 750 um Be TPOL converter from Richard. Alternate w 75um at the end of each set of 4 pol orientation + amo. Will be implemented.